OpenStudy (anonymous):

Find out the values of x which satisfy x^3(x-1)(x-2)^2>0 the answer given is (-infinity,0) U (1,2) U ( 2, infinity) .. please show the steps to solve the problem.

OpenStudy (anonymous):

=taking values of x where function becomes 0 we can find the lowest limit actually :) the numbers axcept 1 and 2 are all satisfying the expression either positive or negitive

OpenStudy (anonymous):

hey x satisfies 2 in this expression first expand and then substitute.....

OpenStudy (anonymous):

no x does not satisfy 2

OpenStudy (anonymous):

if u put x=2 all expression will be 0 and 0>0 is wrong so listion dont putt x=1 or x=2 ..... other real numbers can be ur answers u can check ur self or if u have any number in dobt i can make it clear for u

OpenStudy (lncognlto):

X cannot be 0, 1 or 2, so anything else satisfies the equation. so x is the set of reals, but not equal to 0, 1 or 2.

terenzreignz (terenzreignz):

My favourite way of solving this sort of thing is to take the factors of the expression one by one... But first, I need you to identify what are called the 'critical numbers' of this expression. Critical numbers would be the values of x which would make the left side of your inequality equal to zero...

OpenStudy (rsadhvika):

x=0, x=1, x=2 wud make the left side 0 then...

terenzreignz (terenzreignz):

this isn't your question? @rsadhvika Anyway, we make a table :) \[\Large \left[\begin{matrix}& -\infty & & 0 & & 1 & & 2 & & +\infty\\x^3&&&&&&&&\\(x-1)&&&&&&&& \\(x-2)^2 \\x^3 (x-1)(x-2^2)\end{matrix}\right]\]

OpenStudy (rsadhvika):

yeah i have another method, but since its your fav method, want to see it... .and since the asker is not here, dont want to miss the method :o

terenzreignz (terenzreignz):

this is for beginners, though... :) anyway The critical number associated with \(\large x^3\) \[\Large \left[\begin{matrix}& -\infty & & 0 & & 1 & & 2 & & +\infty\\x^3&&&0&&&&&\\(x-1)&&&|&&&&& \\(x-2)^2&&&|&&&&& \\x^3 (x-1)(x-2^2)\end{matrix}\right]\]

OpenStudy (anonymous):

@rsadhvika you can show your method as well

terenzreignz (terenzreignz):

is zero, sorry.

OpenStudy (rsadhvika):

after this ima post :)

terenzreignz (terenzreignz):

Maybe your method is easier, @rsadhvika ? I wouldn't want to complicate things for @digitalmonk

terenzreignz (terenzreignz):

Anyway, before 0, \(\large x^3 \) takes negative values, and positive values after 0 \[\Large \left[\begin{matrix}& -\infty & & 0 & & 1 & & 2 & & +\infty\\x^3&-&-&0&+&+&+&+&+&+\\(x-1)&&&|&&&&& \\(x-2)^2&&&|&&&&& \\x^3 (x-1)(x-2^2)\end{matrix}\right]\]

OpenStudy (rsadhvika):

oh each factor we will check and apply -x- = + ... i see looks pretty straight forward and cool :)

terenzreignz (terenzreignz):

The critical number associated with (x-1) [the x-value that would make x - 1 = 0] is 1 \[\Large \left[\begin{matrix}& -\infty & & 0 & & 1 & & 2 & & +\infty\\x^3&-&-&0&+&+&+&+&+&+\\(x-1)&&&|&&0&&& \\(x-2)^2&&&|&&|&&& \\x^3 (x-1)(x-2^2)\end{matrix}\right]\] Before 1, (x-1) takes negative values, after 1, x-1 takes positive values... \[\Large \left[\begin{matrix}& -\infty & & 0 & & 1 & & 2 & & +\infty\\x^3&-&-&0&+&+&+&+&+&+\\(x-1)&-&-&-&-&0&+&+&+&+ \\(x-2)^2&&&|&&|&&& \\x^3 (x-1)(x-2^2)\end{matrix}\right]\]

terenzreignz (terenzreignz):

Finally, for \(\large(x-2)^2\) The critical number fot it is 2, but since it's squared, everywhere else, it's going to be positive... \[\Large \left[\begin{matrix}& -\infty & & 0 & & 1 & & 2 & & +\infty\\x^3&-&-&0&+&+&+&+&+&+\\(x-1)&-&-&-&-&0&+&+&+&+ \\(x-2)^2&+&+&+&+&+&+&0&+&+ \\x^3 (x-1)(x-2^2)\end{matrix}\right]\]

terenzreignz (terenzreignz):

Product of the three signs go at the bottom, just remember that - times - is + \[\Large \left[\begin{matrix}& -\infty & & 0 & & 1 & & 2 & & +\infty\\x^3&-&-&0&+&+&+&+&+&+\\(x-1)&-&-&-&-&0&+&+&+&+ \\(x-2)^2&+&+&+&+&+&+&0&+&+ \\x^3 (x-1)(x-2)^2&+&+&0&-&0&+&0&+&+\end{matrix}\right]\] And now just look for the intervals where the prouct is positive, since that's what you needed...

OpenStudy (anonymous):

ok got you @terenzreignz ..thanks a lot..... @rsadhvika please show me your method as well when you get time.

OpenStudy (rsadhvika):

i wud try something like this :- x^3(x-1)(x-2)^2 > 0 Observation 1 : see that its 6 degree polynomial inequality

OpenStudy (rsadhvika):

for even degree polynomials extremes goes in same direction, since leading coefficient is +ve, the extremes goes UP and since it has 3 distinct zeroes, it cuts/touches x axis at 3 places so the domain where it stays UP wud be like this :- (-inf, ) U (, inf)

OpenStudy (rsadhvika):

next visualize its graph :- extremes going UP, adn touching/cuttings x-axis at 3 zeroes : 0, 1, 2

OpenStudy (rsadhvika):